CRU Weekly Meeting



Videoconference Rooms
Erno David
Auto-join URL
Useful links
Phone numbers


Olivier, Tuan, Erno, Tivadar, Joël

Hardware status

(Tivadar, Tuan)

  • TTC-PON testing with 10 CRU did not work.
    • Calibration phase was OK
    • Uplink communication failed. Team diagnosed that we still suffer with the TX disable (not fast enough, because of open collector). We have to go for the push-pull driver, (Vin min for laser is 2.0V while arria 10 output is 2.5V)
    • The CPLD was reverted to the original firmware. Good news no need to replace
    • This afternoon, system level test will be done with the 10 CRU with reprogrammed CRU.
  • PRR answers
    • We have to complete the 10 CRU link testing before sending document


Firmware status

(Joël, Olivier)

  • Rebase to be done after meeting.
  • GBT cdc moving at the output of the transceiver, constraints updated
    • Many re-compilations with small changes, after rebasing. Always working, test done with 24 links firmware (only 16 tested because of the fiber we have)
    • Proposal to provide the firmware to Tuan for testing with the 24 links
    • Now 4h of compilation for the 24 GBT links version (thanks to constraint relaxation) and 2h20 for the 6 links
    • Some of the recompilation have been seen without any timing errors. Now there are some in the PCIe part (please check and correct if possible, either by design or by constraint relaxation)



(Olivier, PiPPo)

  • Still some issues to read out data.
    • Sometimes readout does not work at all → cold or warm reboor fw upload and there it goes, still to be understood
  • LSB/MSB swapping and RDH V3 were implemented, it solved some issues, but not all
  • Still work to do. dwrapper is currently in rewrite to simplify things. In progress
    • postponed many times, we need to release by the end of month with at least the readout working and the cru-test repo up to date.
    • The new wfplayer can arrive a bit later is necessary. All less urgent update will be then continuously released
    • Reminder : MFT test beam by end of june!!


Cru-tests (python)

(Tuan, Olivier)

  • Some cleaning has been done in the repo to use higher level functions and ease script writing. It was used to test the dwrapper
  • Cleaning is still in progress. Tuan now focuses on the TTC pon issues and he'll then come back to the scripts


Power consumption firmware (challenge signal integrity)


  • 4 GBT link, one per bank
  • Olivier will send script instruction for Erno


Trigger meeting


  • Olivier reported its participation to a meeting trigger by ITS in order to establish the "trigger protocol" between FEE and CRU and CRU and CTP.
  • It is different from the regular TTC-PON meeting
  • Discussion is starting by ITS, it will then be widen to all


There are minutes attached to this event. Show them.
The agenda of this meeting is empty
Your browser is out of date!

Update your browser to view this website correctly. Update my browser now